# BICEP/Keck XIX: Extremely Thin Composite Polymer Vacuum Windows for BICEP and Other High Throughput Millimeter Wave Telescopes

### v1

@@ Abstract
1 Introduction
The cosmic microwave background (CMB) is the afterglow from the Big Bang, and as such, is an excellent tool to explore cosmological models (Penzias & Wilson, 1965; Planck Collaboration et al., 2020a). CMB photons—emitted as the opaque primordial plasma cooled into transparent neutral gas—carry information about the conditions of the early universe. The CMB has a near-perfect black body spectrum, ruling out a steady state expanding universe, and the temperature deviations on that spectrum provide strong evidence for the standard model of cosmology, CDM (Wright et al., 1994; Spergel et al., 2003; Planck Collaboration et al., 2020a). There remain unresolved questions about the early universe (such as the horizon and flatness problems) that paradigms like Inflation attempt to explain. Currently, several experiments are aiming to constrain the parameters of Inflation by mapping the CMB. A very small excess of primordial “B modes” (curl in the polarization field of the CMB) at degree scales is an expected prediction from Inflation in the early universe (Kamionkowski & Kovetz, 2016).
@@ 1 Introduction
The entire instrument is placed within a vacuum cryostat to cool both the optics and the sub-kelvin detectors. However, the cryostat obviously requires a transmissive window to observe external signals, which must remain in contact with the ambient environment. Vacuum windows must have low transmission loss in band and be strong enough to withstand the force of atmospheric pressure. Previously, millimeter-wave vacuum windows have been made of ceramics [such as fused-silica (Datta et al., 2021)], plastic foams [including various formulations of Zotefoam (BICEP2 and Keck Array Collaborations et al., 2015)], or bulk plastic materials [primarily high density polyethylene (HDPE) or ultra high molecular weight polyethylene (UHMWPE) (BICEP/Keck Collaboration et al., 2022; D’Alessandro et al., 2018)]. A more detailed history of the development of millimeter vacuum windows can be found in Barkats et al. (2018).
The larger aperture sizes of modern CMB receivers, however, limit the materials available for use. BICEP3 and BICEP Array have nominal clear apertures at the window of 730 mm, while the full outer diameter of the window is 900 mm. Current plans for the CMB-Stage 4 Small Aperture Telescopes (SATs) and Large Aperture Telescope (LATs) include similarly sized windows (Abazajian et al., 2019). Ceramics become prohibitively expensive at those scales. Plastic foams would become impracticably thick and lossy: the primary advantage of the foam windows are their low index of refraction and low transmission loss, but they have the disadvantage of being very weak. A suitably thick foam window would include many laminated layers, which would make their loss properties significantly worse. For example, the 150 GHz Keck receiver windows were 120 mm thick and estimated to cause 2% transmission loss (BICEP2 and Keck Array Collaborations et al., 2015); the doubled aperture size for new BICEP receivers would require a foam window four times thicker. A window with such high transmission loss (8%) would add approximately 3 pW optical load to a 150 GHz detector, swamping out all other sources of optical power at that frequency.
Bulk plastics, being relatively weak though still stronger than plastic foams, must also become significantly thicker to maintain a suitable safety factor for a vacuum vessel (discussed further in Section 2.2). For BICEP3, a 95 GHz receiver, we previously estimated that over half of the instrument loading originated from the 31.8 mm thick HDPE slab window (BICEP/Keck Collaboration et al., 2022). That window has since been replaced by the thin window presented in this paper.
@@
We have achieved an order of magnitude reduction in window thickness using a special form of polyethylene we call high modulus polyethylene (HMPE, commercial name Dyneema (Dyneema, 2008) or Spectra) laminated with low density polyethylene (LDPE) to generated a polyethylene composite material. The high strength of the HMPE allows for thicknesses of around a millimeter, in the same range as the wavelength of light we are observing, which will become relevant for the optical properties of the window. The LDPE melts at lower temperature than the HMPE, allowing for the laminating material to fully surround the HMPE fibers without compromising the strength carrier’s integrity (Vasile, 2005; Eiben et al., 2022). Because the composite is entirely polyethylene, we expect the window to appear homogeneous to millimeter light, and therefore have similar optical properties—such as index of refraction, loss and scattering—as standard bulk polyethylene.
As of 2024, we have deployed two thin laminate windows on BICEP/Keck receivers; one as a retrofit on the 95 GHz BICEP3 and one on the new 150 GHz BICEP Array receiver (hereafter BA150). Section 2 of this paper discusses the design considerations and mechanical tests for these windows. Section 3 explores in-lab validation of optical properties such as polarization and anti-reflection. Section 4 reviews optical characterizations performed on the deployed window before and after installation on BICEP3. We then conclude with final thoughts and next steps.
2 Design and Mechanical Tests
@@
As shown in Figure 2, the window is the first optical element facing the sky. Above the window are “forebaffles” which terminate sidelobe and scattered pickup outside the nominal 27.4∘ field of view of the telescope at a stable, warm temperature (BICEP/Keck Collaboration et al., 2022). Below the window are a stack of physically separated thin polyethylene foam sheets (Zotefoam HD30) which act as an IR filter (Choi et al., 2013). These thin laminate windows must fulfill a variety of criteria to avoid negatively impacting our sensitive instruments. Mechanically, the window should not allow excessive amounts of gas into the cryostat (Section 2.5), must withstand the force of atmospheric pressure with a reasonable safety factor (Sections 2.2 and 2.3), and must not deflect too far inwards to avoid contacting the IR foam filter below them (Section 2.4).
2.1 Manufacturing
These thin windows are composite polyethylene laminates. The HMPE comes in a woven fabric on rolls 1500 mm wide, out of which 900 mm diameter circles are cut with a hot wire; 900 mm is the outer diameter of the frames which hold these windows on the cryostat. At least three layers of HMPE are included in every window, and each layer is rotated 30 degrees from the one below. The number of HMPE layers is primarily set by the maximum allowed deflection and accounts for the expected creep it will experience over its lifetime (more details in Sections 2.2 and 2.4). LDPE layers surround the HMPE: we control the final laminate thickness with the number and initial thickness of LDPE layers. At least four layers of 150 m (6 mil) LDPE are required to fully laminate the three layers of HMPE.
Lamination is processed at high pressure (10 atm) to fully surround the HMPE fibers, which is visually demonstrated in Figure 3; Sections 2.5 and 3.1 both describe mechanical and optical difficulties encountered with poor lamination. The lamination temperature must remain between 130∘C and 140∘C to melt the LDPE around the HMPE without affecting the HMPE’s strength: we typically aim for 135∘C to 137∘C. When appropriate we distinguish between windows by their manufacturing processes, i.e., the pressure and/or temperature of the lamination. The design of the optics autoclave we use for this heat compression process is described further in Eiben et al. (2022).
2.2 Strength
In order to dimension these new thin windows correctly, we first start with the measured mechanical properties: the elastic modulus, the ultimate tensile strength, and the strain at break. We measure these mechanical properties using a tensile strength machine, that collects force versus elongation data over the range of extension of the material up to failure. Those data are then converted into engineering stress (force over the initial cross sectional area) and strain (percent elongation). Example tensile strength measurements for a 10 atm 135∘C HMPE laminate are shown in Figure 4. These measurements are also presented in Table 1 for the HMPE laminate material, along with literature values for the bulk HDPE and UHMWPE found in Barkats et al. (2018); D’Alessandro et al. (2018).
@@ 2.2 Strength
The ratio of the window’s ultimate tensile strength to the predicted stress is the safety factor. We can iterate on the window thickness to achieve a desired safety factor, which we set to be at least three. Section 2.3 presents a detailed motivation for this choice. One way to compare the polyethylene window materials is to calculate the thickness required for each polyethylene window material to reach the required safety factor, using the reported material properties in Table 1 and Equation 2. These minimum safe thicknesses are reported in Table 1; the thickness for a safety factor of three in Table 1 is calculated using the lower value of ultimate tensile strength. Both bulk polyethylenes have similar lower limits on the ultimate tensile strength and therefore require an order of magnitude thicker material to achieve the same safety factor as the HMPE window.
It is important to note that these deflection, stress, and strain calculations use the full thickness of the laminate even though a significant fraction of the thickness is taken up by the much weaker LDPE. As such, the safety factor thicknesses reported in Table 1 are a lower limit. To confirm that the thin laminate windows have an appropriately high margin, we conducted an empirical test of the laminate’s strength, under the clamping boundary conditions that a deployed window would experience, which we discuss in the next section.
2.3 Safety Factor Measurement
Traditional tensile tests like those performed in tensile testing only stress the woven fabric in a single direction (the direction of the pulling force). However, a three layer laminate will have HMPE fibers running in six separate directions over a circular aperture (each layer is rotated 30 degrees from the one below it). A tensile test does not accurately capture the stress conditions of an actual window. The scaling of tensile tests to the full scale conditions thus remains uncertain: the maximum stress and safety factor calculations presented in Section 2.2 rely on the thin plate analytical relationship, which assume homogeneous material, and may not fully apply to composites. To mitigate this concern, we performed a full scale window failure measurement.
@@ 2.3 Safety Factor Measurement
We reached 68 psi in the hydrostatic test using only the standard utility water supply. At that point we took a side-on picture of the deflected window, and estimated the deflection (and therefore radial strain) of the window (see Section 2.4 for a description of the calculations). The window had deflected approximately 75 mm. which corresponds to a radial strain of approximately 3%, well below the typical failure strain of 10%. In an attempt to produce a failure, we added compressed air to the system, increasing the pressure to 85 psi. At this pressure water began to leak around the gaskets, and the test was stopped. The end state of the window is shown in Figure 5, with the water pressure valve showing the final pressure on the right and the water leaking out around the gaskets on the left. We note that 85 psi is twice the predicted failure pressure for a safety factor of 3, a degree of uncertainty that emphasizes the importance of empirical testing of this novel structure.
Given these results, we find that the safety factor of the thin windows is at least approximately 5.7 at sea level (1 atm) and 8.2 at South Pole (0.7 atm). Because the window will never experience a pressure over one atmosphere, we can be confident that the window will not fail under normal conditions. However, we must also explore other time-dependent factors which affect the mechanical properties of these window materials.
2.4 Creep
@@
where is the initial radius and is the center deflection. We can relate the estimated radial strain to a power law similar to that in Findley & Tracy (1974), obtained from observations of creep in polyethylene over sixteen years.
Figure 6 shows the measured deflection, fit deflection, and radial strain on a thin window on the PreSAT cryostat over approximately a month (Petroff et al., 2024). The radial strain is predicted to reach 4.5% after 5 years of operation, well below the failure strains of 9–12% (reported in Section 2.2). The window also remains out of contact with the foam filters below it.
2.5 Gas Permeation
@@
These rates can also be compared to other measured atmospheric permeation rates for HDPE, such as those found in Nguyen‐Tuong (1993). At a given temperature, the permeation rate depends on the permeability of a material , the pressure difference between the two volumes (for our purposes assumed to be atmospheric pressure at a given site), and the thickness ():
@@
In 2018, a faulty o-ring seal at the window flange allowed gas to enter the cryostat over the course of one year. The gas accumulation was visible as temperature spikes on the 50 kelvin shell. During the following summer season we warmed the receiver and measured the pressure inside the cryostat. The final pressure was 49 mbar after 359 days under vacuum. If all of that gas had instead permeated through the window, the window would have had a permeation rate of 3.4 mbar L s-1 mm-2. The measured bolometer white noise did not increase significantly in 2018 compared to 2017, implying that this amount of gas accumulation did not adversely impact performance (see Section 4.3 for further discussion on BICEP3’s white noise performance).
In 2022 we saw temperature variations again thought to be due to gas entry into the cryostat. These leaks were localized to a flange at the base of the cryostat. This rate of gas entry resulted in a significant increase in the measured white noise (again, see Section 4.3 for further discussion) for approximately 463 days. However, we cannot distinguish what fraction of the gas entered during the higher noise period and what fraction entered in the previous 626 days the receiver was under vacuum. Conservatively, we average the final warm pressure of 314 mbar over the entire period the receiver was under vacuum, recognizing that it is likely that the actual problematic permeation rate was higher than this average, and posit that a permeation rate through a thin window of 7.2 mbar L s-1 mm-2 will cause significant cryogenic and optical issues. For both the 2018 and 2022 cases, time at vacuum, final warm pressure and estimated upper bound window permeation rate are summarized in the bottom of Table 2.
@@
To determine the actual laminate windows’ permeation rate we tested two small scale samples (shown in the top of Figure 3) on a small vacuum vessel over a few days. Previously, we had noted an unacceptably high permeation rate in thin windows that were laminated at one atmosphere of pressure (Barkats et al., 2018; Eiben et al., 2022). Therefore, we also compared the permeation rates of low pressure laminates and high pressure laminates, to confirm that higher pressure lamination improved the permeation rate. Figure 7 shows the results of the small scale permeation rate testing. The measured small scale permeation rates are compared to the predicted permeation rate for bulk HDPE of equivalent thickness (1.4 mm) using Equation 5 with the permeation constant found in Nguyen‐Tuong (1993).
@@
We have since run an observing season with a high pressure laminate window on BA150 at the South Pole (at altitude, with assumed 0.7 atm pressure difference) and warmed the receiver. We also measured the gas accumulation in another receiver (PreSAT) at sea level, though the run was over a shorter period of time (Petroff et al., 2024). Table 2 reports the time at vacuum and final pressures for each run. The final warm pressure can be used to estimate an upper bound permeation rate through a thin window, which can be compared to the predicted permeation rate for a given window, or the upper bound window permeation rates associated with cryogenic issues.
Both of the receivers with thin windows cooled and remained cold successfully. These measured amounts of gas are also significantly below the previous problematic gas amounts. Based on the experiences described here, CMB-S4 has taken a requirement of 1 mbar/month for similarly scaled cryostats; these window permeation rates will easily achieve this requirement. Roughly, we may estimate that the thin windows may allow a cryogenically problematic amount of gas into the cryostat after about twenty years, wherein the gas could be removed by warming, pumping out, and cooling again. The BICEP/Keck collaboration has very rarely kept a cryostat cold for longer than three consecutive years; permeation of gas through the window into the cryostat is unlikely to become a problem.
3 Optical Tests
The vacuum window is the single highest contributor to optical noise in the instrument. As such, characterizing and minimizing the optical impact of the window on our instrument is critical. After mechanical considerations are satisfied, we design the window to minimize reflected power across our observing band (Section 3.1), and characterize the potential polarizing effect of the laminates at our frequencies (Section 3.2). We also report in-situ characterization on the BICEP3 telescope, via direct optical measurements between different window materials with the CMB detectors (Section 4).
3.1 Anti-reflection

### v2

@@ Abstract
I Introduction
The cosmic microwave background (CMB) is the afterglow from the Big Bang, and as such, is an excellent tool to explore cosmological models (Penzias and Wilson, 1965; Planck Collaboration et al., 2020a). CMB photons—emitted as the opaque primordial plasma cooled into transparent neutral gas—carry information about the conditions of the early universe. The CMB has a near-perfect black body spectrum, ruling out a steady state expanding universe, and the temperature deviations on that spectrum provide strong evidence for the standard model of cosmology, CDM (Wright et al., 1994; Spergel et al., 2003; Planck Collaboration et al., 2020a). There remain unresolved questions about the early universe (such as the horizon and flatness problems) that paradigms like inflation attempt to explain. Currently, several experiments are aiming to constrain the parameters of inflation by mapping the CMB. A very small excess of primordial “B modes” (curl in the polarization field of the CMB) at degree scales is an expected prediction from Inflation in the early universe (Kamionkowski and Kovetz, 2016).
@@ I Introduction
The entire instrument is placed within a vacuum cryostat to cool both the optics and the sub-kelvin detectors. However, the cryostat requires a transmissive window to observe external signals, which must remain in contact with the ambient environment. Vacuum windows must have low transmission loss in band and be strong enough to withstand the force of atmospheric pressure. Previously, millimeter-wave vacuum windows have been made of ceramics [such as fused-silica (Datta et al., 2021)], plastic foams [including various formulations of Zotefoam (BICEP2 and Keck Array Collaborations et al., 2015)], or bulk plastic materials [primarily high density polyethylene (HDPE) or ultra high molecular weight polyethylene (UHMWPE) (BICEP/Keck Collaboration et al., 2022; D’Alessandro et al., 2018)]. A more detailed history of the development of millimeter vacuum windows can be found in Barkats et al. (2018).
The larger aperture sizes of modern CMB receivers, however, limit the materials available for use. BICEP3 and BICEP Array have nominal clear apertures at the window of 730 mm, while the full outer diameter of the window is 900 mm. Plans for the CMB-S4 Small Aperture Telescopes (SATs) and Large Aperture Telescope (LATs) included similarly sized windows (Abazajian et al., 2019). Ceramics become prohibitively expensive at those scales. Plastic foams would become impracticably thick and lossy: the primary advantages of the foam windows are their low index of refraction and low transmission loss, but they have the disadvantage of being very weak. A suitably thick foam window would include many laminated layers, which would make their loss properties significantly worse. For example, the 150 GHz Keck receiver windows were 120 mm thick and estimated to cause 2% transmission loss (BICEP2 and Keck Array Collaborations et al., 2015); the doubled aperture size for new BICEP receivers would require a foam window four times thicker. A window with such high transmission loss (8%) would add approximately 3 pW optical load to a 150 GHz detector, swamping out all other sources of optical power at that frequency.
Bulk plastics, being relatively weak though still stronger than plastic foams, must also become significantly thicker to maintain a suitable safety factor for a vacuum vessel (discussed further in Section II.2). For BICEP3, a 95 GHz receiver, we previously estimated that over half of the instrument loading originated from the 31.8 mm thick HDPE slab window (BICEP/Keck Collaboration et al., 2022). That window has since been replaced by the thin window presented in this paper.
@@
We have achieved an order of magnitude reduction in window thickness using a special form of polyethylene we call high modulus polyethylene (HMPE, commercial name Dyneema (Dyneema, 2008) or Spectra) laminated with low density polyethylene (LDPE) to generated a polyethylene composite material.
As of 2025, we have deployed three thin laminate windows on BICEP/Keck receivers; one as a retrofit on the 95 GHz BICEP3, one on the new 150 GHz BICEP Array receiver (hereafter BA150) and another on the new 220/270 GHz BICEP Array receiver (hereafter BA220/270). Section II of this paper discusses the design considerations and mechanical tests for these windows. Section III explores in-lab validation of optical properties such as polarization and anti-reflection. Section IV reviews optical characterizations performed on the deployed window before and after installation on BICEP3. We then conclude with final thoughts and next steps.
II Design and Mechanical Tests
The high strength of the HMPE allows for window thicknesses of around a millimeter—similar to the wavelength of light we are observing, which will become relevant for the optical properties of the window. The LDPE melts at lower temperature than the HMPE, allowing for the laminating material to fully surround the HMPE fibers without compromising the strength carrier’s integrity (Vasile, 2005; Eiben et al., 2022). Because the composite is entirely polyethylene, one should expect the window to appear homogeneous to millimeter light, and therefore have similar optical properties—such as index of refraction, loss and scattering—as standard bulk polyethylene.
@@ II Design and Mechanical Tests
As shown in Figure 2, the window is the first optical element facing the sky. Above the window are “forebaffles” which terminate sidelobe and scattered pickup outside the nominal 27.4∘ field of view of the telescope at a stable, warm temperature (BICEP/Keck Collaboration et al., 2022). Below the window are a stack of physically separated thin polyethylene foam sheets (Zotefoam HD30) which act as an IR filter (Choi et al., 2013; Goldfinger et al., 2022). These thin laminate windows must fulfill a variety of criteria to avoid negatively impacting our sensitive instruments. Mechanically, the window should not allow excessive amounts of gas into the cryostat (Section II.5), must withstand the force of atmospheric pressure with a reasonable safety factor (Sections II.2 and II.3), and must not deflect too far inwards to avoid contacting the IR foam filter below them (Section II.4).
II.1 Manufacturing
The thin windows are composite polyethylene laminates. The HMPE comes in a woven fabric on rolls 1500 mm wide, out of which 900 mm diameter circles are cut with a hot wire; 900 mm is the outer diameter of the frames which hold these windows on the cryostat. At least three layers of HMPE are included in every window, and each layer is rotated 30 degrees from the one below. The number of HMPE layers is primarily set by the maximum allowed deflection and accounts for the expected creep it will experience over its lifetime (more details in Sections II.2 and II.4). LDPE layers surround the HMPE: we control the final laminate thickness with the number and initial thickness of LDPE layers. At least four layers of 100 m (4 mil) LDPE are required to fully laminate the three layers of HMPE.
Lamination is processed at high pressure (10 atm) to fully surround the HMPE fibers, which is visually demonstrated in Figure 3; Sections II.5 and III.1 both describe mechanical and optical difficulties encountered with poor lamination. We determined that lamination above 4 atm was necessary for homogeneity, and that higher pressure lamination was more likely to keep the appropriate laminating force during a cycle. The highest pressure we can achieve is 10 atm, so all laminates were baked at 10 atm unless otherwise indicated. The lamination temperature must remain between 130∘C and 140∘C to melt the LDPE around the HMPE without affecting the HMPE’s strength: we typically aim for 135∘C to 137∘C. When appropriate we distinguish between windows by their manufacturing processes, i.e., the pressure and/or temperature of the lamination. The design of the optics autoclave we use for this heat compression process is described further in Eiben et al. (2022).
II.2 Strength
@@
In order to dimension these new thin windows correctly, we first start with the measured mechanical properties: the elastic modulus, the ultimate tensile strength, and the strain at break. We measure these mechanical properties using a tensile strength machine that collects force versus elongation data over the range of extension of the material up to failure. Those data are then converted into engineering stress (force over the initial cross sectional area) and strain (percent elongation). Example tensile strength measurements for a 10 atm 135∘C HMPE laminate are shown in Figure 4. These measurements are also presented in Table 1 for the HMPE laminate material, along with literature values for the bulk HDPE and UHMWPE found in Barkats et al. (2018); D’Alessandro et al. (2018).
The ratio of the window’s ultimate tensile strength to the predicted stress is the safety factor. We can iterate on the window thickness to achieve a desired safety factor, which we set to be at least three. Section II.3 presents a detailed motivation for this choice. One way to compare the polyethylene window materials is to calculate the thickness required for each polyethylene window material to reach the required safety factor, using the reported material properties in Table 1 and Equation 2. We use the lower value in the range of reported ultimate tensile strengths for a given material to determine the minimum safe thickness reported in Table 1. Both bulk polyethylenes have similar lower limits on the ultimate tensile strength and therefore require an order of magnitude thicker material to achieve the same safety factor as the HMPE window.
It is important to note that these deflection, stress, and strain calculations use the full thickness of the laminate even though a significant fraction of the thickness is taken up by the much weaker LDPE. As such, the safety factor thicknesses reported in Table 1 are a lower limit. Previously, we reported the strength of laminates using the HMPE fiber cross section (Barkats et al., 2018). We find that recent laminates have the same high ultimate tensile strength (at or above 3 GPa) when normalizing tensile results to the fiber cross section, which indicates that the lamination process does not reduce the HMPE’s strength. To confirm that the thin laminate windows have an appropriately high margin to their ultimate strength, we conducted an empirical test of the laminate’s strength, under the clamping boundary conditions that a deployed window would experience, which we discuss in the next section.
II.3 Safety Factor Measurement
Traditional tensile tests like those shown in Figure 4 will only stress the woven fabric in a laminate in a single direction (the direction of the pulling force). However, a three layer laminate will have HMPE fibers running in six separate directions over a circular aperture (each layer is rotated 30 degrees from the one below it). A tensile test does not accurately capture the stress conditions of an actual window. The scaling of tensile tests to the full scale conditions thus remains uncertain: the maximum stress and safety factor calculations presented in Section II.2 rely on the thin plate analytical relationship, which assume homogeneous material, and may not fully apply to composites. To mitigate this concern, we performed a full scale window failure measurement.
@@ II.3 Safety Factor Measurement
We reached 68 psi (469 kPa) in the hydrostatic test using only the standard utility water supply. At that point we took a side-on picture of the deflected window, and estimated the deflection (and therefore radial strain) of the window (see Section II.4 for a description of the calculations). The window had deflected approximately 75 mm. which corresponds to a radial strain of approximately 3%, well below the typical failure strain of 10%. In an attempt to produce a failure, we added compressed air to the system, increasing the pressure to 85 psi. At this pressure water began to leak around the gaskets, and the test was stopped. The end state of the window is shown in Figure 5, with the water pressure valve showing the final pressure on the right and the water leaking out around the gaskets on the left. We note that 85 psi (586 kPa) is twice the predicted failure pressure for a safety factor of 3, a degree of uncertainty that emphasizes the importance of empirical testing of this novel structure.
Given these results, we find that the lower limit of the safety factor of the thin windows is approximately 5.7 at sea level (1 atm) and 8.2 at South Pole (0.7 atm). Because the window will never experience a pressure over one atmosphere, we can be confident that the window will not fail under normal conditions. However, we must also explore other time-dependent factors which affect the mechanical properties of these window materials.
II.4 Creep
@@
where is the initial radius and is the center deflection. We can relate the estimated radial strain to a power law similar to that in Findley and Tracy (1974), obtained from observations of creep in polyethylene over sixteen years.
Figure 6 shows the measured deflection, fit deflection, and radial strain on a thin window on the BA4 cryostat over approximately a month (Petroff et al., 2024). The radial strain is predicted to reach 4.5% after 5 years of operation, well below the failure strains of 9–12% (reported in Section II.2). The window also remains out of contact with the foam filters below it.
II.5 Gas Permeation
@@
These rates can also be compared to other measured atmospheric permeation rates for HDPE, such as those found in Nguyen‐Tuong (1993). At a given temperature, the permeation rate depends on the permeability of a material , the pressure difference between the two volumes (for our purposes assumed to be atmospheric pressure at a given site), the thickness () and the area over which permeation occurs ():
@@
In 2018, a faulty o-ring seal at the window flange allowed gas to enter the cryostat over the course of one year. During the following summer season we warmed the receiver and measured the pressure inside the cryostat. The final pressure was 49 mbar after 359 days under vacuum. If all of that gas had instead permeated through the window, the window would have had a permeation rate of 3.4 mbar L s-1 mm-2. The measured bolometer white noise did not increase significantly in 2018 compared to 2017, implying that this amount of gas accumulation did not adversely impact performance (see Section IV.3 for further discussion on BICEP3’s white noise performance).
In 2022 leaks in a flange at the base again allowed excess gas entry into the cryostat. This rate of gas entry resulted in a significant increase in the measured white noise (again, see Section IV.3 for further discussion) for approximately 463 days. However, we cannot distinguish what fraction of the gas entered during the higher noise period and what fraction entered in the previous 626 days the receiver was under vacuum. Conservatively, we average the final warm pressure of 314 mbar over the entire period the receiver was under vacuum, recognizing that it is likely that the actual problematic permeation rate was higher than this average, and posit that a permeation rate through a thin window of 7.2 mbar L s-1 mm-2 will cause significant cryogenic and optical issues. For both the 2018 and 2022 cases, time at vacuum, final warm pressure and estimated upper bound window permeation rate are summarized in the bottom of Table 2.
@@
To isolate the laminate windows’ permeation rate from other sources of gas accumulation in a vacuum, we tested two small scale samples (shown in the top of Figure 3) on a small vacuum vessel over a few days. Previously, we had noted an unacceptably high permeation rate in thin windows that were laminated at one atmosphere of pressure (Barkats et al., 2018; Eiben et al., 2022). Therefore, we also compared the permeation rates of low pressure laminates and high pressure laminates, to confirm that higher pressure lamination reduced the permeation rate. Figure 7 shows the results of the small scale permeation rate testing. The measured small scale permeation rates are compared to the predicted permeation rate for bulk HDPE of equivalent thickness (1.4 mm) using Equation 5 with the permeation constant found in Nguyen‐Tuong (1993).
@@
We have since run an observing season with a high pressure laminate window on a cryostat (BA150) at the South Pole (at altitude, with assumed 0.7 atm pressure difference) and warmed the receiver. We also measured the gas accumulation in another receiver (BA4) at sea level, though the run was over a shorter period of time (Petroff et al., 2024). Table 2 reports the time at vacuum and final pressures for each run. The final warm pressure can be used to estimate an upper bound permeation rate through a thin window, which can be compared to the predicted permeation rate for a given window, or the upper bound window permeation rates associated with cryogenic issues.
Both of the receivers with thin windows cooled and remained cold successfully. The predicted window permeation rates are below the upper bound rate estimated from the total gas accumulation, indicating that the permeation through the windows is a significant contributor to gas accumulation inside the cryostat. The window likely dominates over other sources of gas accumulation (such as permeation through the o-rings) over long timescales. These measured amounts of gas are also significantly below the previous problematic gas amounts. Roughly, we may estimate that the thin windows may allow a cryogenically problematic amount of gas into the cryostat after about twenty years, wherein the gas could be removed by warming, pumping out, and cooling again. The BICEP/Keck collaboration has very rarely kept a cryostat cold for longer than three consecutive years; permeation of gas through the window into the cryostat is unlikely to become a problem.
III Optical Tests
The vacuum window is the single highest contributor to optical noise in the instrument. As such, characterizing and minimizing the optical impact of the window on our instrument is critical. After mechanical considerations are satisfied, we design the window to minimize reflected power across our observing band (Section III.1), and characterize the potential polarizing effect of the laminates at our frequencies (Section III.2). We also report in-situ characterization on the BICEP3 telescope, via direct optical measurements between different window materials with the CMB detectors (Section IV).
III.1 Anti-reflection